# Controlling Multimodal Conversational Agents with Coverage-Enhanced Latent Actions

### v1

@@ 5 Related Work
Recent advances in vision-language models (VLMs) bai-2025-qwen25vl have enabled increasingly capable multimodal conversational agents (MCAs) yao-2025-MLLMAgentsurvey, such as multimodal role-playing agents dai-2025-mmrole and personalized assistants nguyen-2024-yo; li-2025-aligning, which hold significant promise in fields like entertainment mehta-2022-exploring and personalized education griol-2014-developing. Initial efforts to build MCAs primarily rely on supervised fine-tuning lillava-2024-TMLR, but often suffer from poor generalization. Recently, RL has been widely explored for fine-tuning MCAs and has demonstrated strong generalization performance zhou-2025-reinforcedMLLM; chu-2025-sftRL. However, fine-tuning MCAs via RL faces challenges in handling the extremely large text token space. To address this, we propose constructing a compact latent action space for RL fine-tuning, which enables efficient policy learning.
@@
In many real-world scenarios, only observation-only data are available, such as expert demonstration videos of robots where explicit action labels are missing Torabi-2019-RecentImitation. To address this challenge, prior works leverage the learning from observation mechanism seo-2022-reinforcement; baker-2022-video to infer latent actions from observation-only data, which are then used for RL fine-tuning of agents. For instance, zhang-2024-whale; gao-2025-adaworld learn latent actions from videos to control video generation, while ye-2025-latent; bu-2025-univla extract latent actions from robot manipulation videos and use them for robot policy learning. These constructed latent actions not only enhance controllability bruce-2024-genie but also enable better transferability across different tasks due to their higher-level nature jang-2025-dreamgen.
The most relevant work to ours is CoLA jia-2025-controlling, which introduces latent actions into RL fine-tuning of LLMs. However, when constructing the latent action space for multimodal conversational agents, the scarcity of paired image-text data hinders learning a latent space with sufficient coverage. To overcome this, we leverage both paired image-text data and massive text-only data to construct the latent space, using a cross-modal projector trained with a novel cycle-consistency loss.
@@ 6 Conclusion
We acknowledge the following limitations in our work. First, the additional latent action prediction step increases RL training time by 1.08× and inference latency by 1.13×. Second, due to constraints of computational resources, we evaluate our latent action based approach on multimodal conversational tasks, and leave validation on more tasks to future work, such as visual mathematical reasoning.
@@
Our work is entirely at the methodological level, which means that there will not be any negative social impacts.
@@ A.2 Inverse Dynamics Model
Following jia-2025-controlling, we adopt a direct code assignment strategy to avoid code collapse. Specifically, a linear head (inverse action head) maps $\tilde{e}^{V,T}_{t+1}$ to logits $\mathbf{l}_{t}\in\mathbb{R}^{|\mathcal{C}|}$ over the codebook indices. During inverse dynamics learning, we apply the Gumbel-Softmax and a reparameterization trick to obtain a differentiable soft assignment:
@@ A.4 Codebook for the Latent Action Space
The latent action space is defined by a codebook $\mathcal{C}=\{c_{1},\dots,c_{K}\}\subset\mathbb{R}^{d}$ with $K=128$. Each code vector $c_{k}$ is initialized independently via Kaiming uniform initialization he-2015-delving. Given a latent action index $a_{t}\in\{1,\dots,K\}$, the corresponding latent action embedding is retrieved as $c_{a_{t}}\in\mathcal{C}$.
@@ B.1 Details on Datasets
To construct the latent action space in an unsupervised manner, we collect large-scale paired image-text and text-only corpora. For paired image-text data, we use: (1) image-caption pairs from Conceptual-12M changpinyo-2021-Conceptual; (2) multimodal news articles from N24News wang-2022-N24News; and (3) multimodal Wikipedia articles from WikiWeb2M burns-2023-wiki, comprising 14M images and 1B text tokens in total. For text-only data, we primarily sample 500K sequences from SlimPajama-627B cerebras-2023-slimpajama due to computational constraints, and additionally include 40K alignment corpora from HelpSteer3 wang-2025-helpsteer3 to preserve the original VLM’s safety and preference alignment during latent space learning. To ensure fair comparison, we analyze data exposure in Appendix C.1 and find that downstream task performance does not benefit from the above corpora, confirming that observed improvements stem from methodological advances.
@@ B.2 Details on Evaluation Metric
We adopt LLM-as-a-Judge metrics to evaluate model performance, using prompt templates validated by dai-2025-mmrole; li-2025-aligning, which show high correlation with human judgments. The evaluation prompt templates used on MMRole and PCogAlignBench are shown in Table 4. We adopt the Qwen3-235B-A22B by the Qwen3 API platform as the judge model.
@@ B.3 Training Details
All experiments are conducted on a single machine equipped with 4 Nvidia A100-80G GPU. For the baseline SFT and RL algorithms, as well as our newly proposed latent action RL methods, we adapt the framework based on the TRL library vonwerra-2022-trl.
@@ B.4 Inference Details
For all methods, we use a sampling temperature of 0.1 during inference, i.e., for token-based baselines, this temperature is applied to the token logits; for our latent action based methods, it is applied to the latent action logits. Additionally, following jia-2025-controlling, for our latent action based methods, token generation by the language world model is deterministic, i.e., tokens are selected via argmax over the output token logits.
@@
Table 4: Prompt templates used for LLM-as-a-Judge evaluation on MMRole and PCogAlignBench. These templates follow established designs from dai-2025-mmrole; li-2025-aligning and have been shown to achieve high correlation with human judgments.
@@
Appendix C Additional Empirical Results
C.1 Analysis on Data Exposure
C.2 Detailed Results on Fine-grained Dimensions
We report the fine-grained performance across each evaluation dimensions, previously summarized in Fig. 4. Specifically, Tables 6 and 7 present results on the in-distribution (ID) and out-of-distribution (OOD) splits of MMRole, respectively. Tables 8 and 9 show results on the LS1 and LS2 subsets of PCogAlignBench. All results are obtained using the Qwen2.5-VL-3B-Instruct model.
C.3 Case Study

### v2

@@ 5 Related Work
Recent advances in vision-language models (VLMs) Bai et al. (2025b) have enabled increasingly capable multimodal conversational agents (MCAs) Yao et al. (2025), such as multimodal role-playing agents Dai et al. (2025) and personalized assistants Nguyen et al. (2024); Li et al. (2025a), which hold significant promise in fields like entertainment Mehta et al. (2022) and personalized education Griol et al. (2014). Initial efforts to build MCAs primarily rely on supervised fine-tuning Li et al. (2024), but often suffer from poor generalization. Recently, RL has been widely explored for fine-tuning MCAs and has demonstrated strong generalization performance Zhou et al. (2025); Chu et al. (2025). However, fine-tuning MCAs via RL faces challenges in handling the extremely large text token space. To address this, we propose constructing a compact latent action space for RL fine-tuning, which enables efficient policy learning.
@@
In many real-world scenarios, only observation-only data are available, such as expert demonstration videos of robots where explicit action labels are missing Torabi et al. (2019). To address this challenge, prior works leverage the learning from observation mechanism Seo et al. (2022); Baker et al. (2022) to infer latent actions from observation-only data, which are then used for RL fine-tuning of agents. For instance, Zhang et al. (2024b); Gao et al. (2025) learn latent actions from videos to control video generation, while Ye et al. (2025); Bu et al. (2025) extract latent actions from robot manipulation videos and use them for robot policy learning. These constructed latent actions not only enhance controllability Bruce et al. (2024) but also enable better transferability across different tasks due to their higher-level nature Jang et al. (2025).
The most relevant work to ours is CoLA Jia et al. (2025), which introduces latent actions into RL fine-tuning of LLMs. However, when constructing the latent action space for multimodal conversational agents, the scarcity of paired image-text data hinders learning a latent space with sufficient coverage. To overcome this, we leverage both paired image-text data and massive text-only data to construct the latent space, using a cross-modal projector trained with a novel cycle-consistency loss.
@@ 6 Conclusion
[…]
@@
While our work is primarily methodological, we acknowledge potential ethical concerns inherent in the benchmarks we utilize. Specifically, as noted by PCogAlignBench Li et al. (2025a), role-set bias may arise from incomplete data collection, where user personas used for evaluation may not fully represent diverse real-world scenarios and could inadvertently reflect societal stereotypes. Although such limitations are acceptable in a controlled research environment, we encourage industry practitioners to consider diverse user backgrounds when constructing user personas for data collection and personalized alignment training, in order to mitigate the risk of biases in real-world deployments.
Acknowledgments
This work was supported by the grant from the National Natural Science Foundation of China (NSFC) project (No. 62576256), and the Fundamental Research Funds for the Central Universities, China (Grant No. 2042022dx0001).
@@ A.2 Inverse Dynamics Model
Following Jia et al. (2025), we adopt a direct code assignment strategy to avoid code collapse. Specifically, a linear head (inverse action head) maps $\tilde{e}^{V,T}_{t+1}$ to logits $\mathbf{l}_{t}\in\mathbb{R}^{|\mathcal{C}|}$ over the codebook indices. During inverse dynamics learning, we apply the Gumbel-Softmax and a reparameterization trick to obtain a differentiable soft assignment:
@@ A.4 Codebook for the Latent Action Space
The latent action space is defined by a codebook $\mathcal{C}=\{c_{1},\dots,c_{K}\}\subset\mathbb{R}^{d}$ with $K=128$. Each code vector $c_{k}$ is initialized independently via Kaiming uniform initialization He et al. (2015). Given a latent action index $a_{t}\in\{1,\dots,K\}$, the corresponding latent action embedding is retrieved as $c_{a_{t}}\in\mathcal{C}$.
@@ B.1 Details on Datasets
To construct the latent action space in an unsupervised manner, we collect large-scale paired image-text and text-only corpora. For paired image-text data, we use: (1) image-caption pairs from Conceptual-12M Changpinyo et al. (2021); (2) multimodal news articles from N24News Wang et al. (2022); and (3) multimodal Wikipedia articles from WikiWeb2M Burns et al. (2023), comprising 14M images and 1B text tokens in total. For text-only data, we primarily sample 500K sequences from SlimPajama-627B Soboleva et al. (2023) due to computational constraints, and additionally include 40K alignment corpora from HelpSteer3 Wang et al. (2025) to preserve the original VLM’s safety and preference alignment during latent space learning. To ensure fair comparison, we analyze data exposure in Appendix C.2 and find that downstream task performance does not benefit from the above corpora, confirming that observed improvements stem from methodological advances.
@@ B.2 Details on Evaluation Metric
We adopt LLM-as-a-Judge metrics to evaluate model performance, using prompt templates validated by Dai et al. (2025); Li et al. (2025a), which show high correlation with human judgments. The evaluation prompt templates used on MMRole and PCogAlignBench are shown in Table 4. We adopt the Qwen3-235B-A22B by the Qwen3 API platform as the judge model.
@@ B.3 Training Details
All experiments are conducted on a single machine equipped with 4 Nvidia A100-80G GPU. For the baseline SFT and RL algorithms, as well as our newly proposed latent action RL methods, we adapt the framework based on the TRL library von Werra et al. (2020).
@@ B.4 Inference Details
For all methods, we use a sampling temperature of 0.1 during inference, i.e., for token-based baselines, this temperature is applied to the token logits; for our latent action based methods, it is applied to the latent action logits. Additionally, following Jia et al. (2025), for our latent action based methods, token generation by the language world model is deterministic, i.e., tokens are selected via argmax over the output token logits.
@@
Table 4: Prompt templates used for LLM-as-a-Judge evaluation on MMRole and PCogAlignBench. These templates follow established designs from Dai et al. (2025); Li et al. (2025a) and have been shown to achieve high correlation with human judgments.
Appendix C Additional Empirical Results
C.1 Analysis on Codebook Size
To investigate the effect of codebook size, we ablate over three settings: 64, 128, and 256. As shown in Table 5, all three sizes yield comparable performance, indicating that our method is robust to the choice of codebook size. We adopt 128 as the default setting.
Codebook Size MMRole PCogAlignBench Avg.
ID OOD LS1 LS2
64 0.946±0.005 0.914±0.066 0.875±0.009 0.848±0.011 0.896±0.023
128 0.949±0.007 0.915±0.065 0.871±0.011 0.837±0.010 0.893±0.023
256 0.953±0.008 0.921±0.032 0.874±0.010 0.838±0.008 0.897±0.015
Table 5: Analysis on codebook size. We evaluate on MMRole and PCogAlignBench using the LLM-as-a-Judge metric. Results are averaged over three runs. All variants are fine-tuned with GRPO (Latent Action) based on Qwen2.5-VL-3B-Instruct.
C.2 Analysis on Data Exposure
@@
C.3 Detailed Results on Fine-grained Dimensions
We report the fine-grained performance across each evaluation dimensions, previously summarized in Fig. 4. Specifically, Tables 7 and 8 present results on the in-distribution (ID) and out-of-distribution (OOD) splits of MMRole, respectively. Tables 9 and 10 show results on the LS1 and LS2 subsets of PCogAlignBench. All results are obtained using the Qwen2.5-VL-3B-Instruct model.
C.4 Case Study